# These Are Not All the Features You Are Looking For: A Fundamental Bottleneck in Supervised Pretraining

### v2

@@ 4.2 Spurious Features
Size Base-Pairs Score $\uparrow$ Rank $\downarrow$ %Imp.$\uparrow$ %Imp.$\uparrow$
@@
4.3 Genomic Foundation Models
@@
4.4 Rich Representations: The Beginnings of a Solution?
(Zhang and Bottou, 2023) propose learning richer representations to capture these missing features. They construct models by concatenating feature representations of individual models and demonstrate significant improvements in accuracy of transfer.
@@ 4.4 Rich Representations: The Beginnings of a Solution?
Table 2: Supervised transfer learning from ImageNet to iNat18, Cifar100, and Cifar10 using linear probing. The cat4 ensemble concatenates representations of $4$ separately trained networks. The baseline uses a ResNet model of similar size is trained once. These models contain approximately four times the parameters of the original ResNet50.
@@
ensemble cat4$\times$×resnet50 94m 78.15 46.55 78.19 93.09
They find the concatenated ensemble to perform comparably to the baseline on ImageNet1K, but far better when transferred to new datasets, ($+9\%$ on INat18) and ($+5.5\%$ on CIFAR100). Additional comparison to the original ResNet50 reveals that only increasing size of a model (such as the baseline) will worsen performance of transfer: (.$-.5\%$ on INat18), and (.$-.6\%$ on CIFAR100).
@@
This also explains why performance on ImageNet1K remains relatively unchanged. Observe that these subsets of features work similarly on ImageNet, that is, each set of features provides the same information to reduce training error. Combining these features marginally improves performance on ImageNet1K ($+2\%$). Hence, they are not learned during training.
@@
4.5 Connections to Multi-Calibration
Missing features fundamentally limits the ability of a model to generalize to new distributions outside the training distribution (OOD). (Wald et al., 2021) finds a link between OOD performance and multi-calibration, introduced by (Hebert-Johnson et al., 2018). Specifically, they argue that calibration leads to better OOD generalization, by creating invariant representations (Arjovsky et al., 2019). (Guy et al., 2025) successfully demonstrate this empirically for in-distribution generalization (IID). However, any remediation requires recovering these missing features.
5 Experiments
5.1 Rich Representations That Cost Nothing
We propose an analogous method to construct the rich representations, inspired by (Zhang et al., 2022), which concatenates networks pretrained for fewer steps. Unlike previous cat experiments that preserve model size (i.e. - space), our method preserves pretraining compute (time). Our findings reveal that the (time-cat) method improves generalization to new datasets while maintaining the cost of pretraining.
5.1.1 Method
Our approach concatenates feature representations from multiple ResNet50 pretrained on ImageNet1k for a shorter duration (Deng et al., 2009), (He et al., 2016).  [^4] Implementation can be found here: https://github.com/facebookresearch/richreps-timecat Each ResNet50 networks is trained identically except using different random seeds. We transfer these concatenated networks using linear probing on new datasets: iNaturalist18, CIFAR-100, CIFAR-10 (Van Horn et al., 2018), (Krizhevsky, 2009).
Our baseline uses a single ResNet50 model trained on ImageNet1k for 4̃50k steps, (reaching 0.65 accuracy on ImageNetv2). We compare how this baseline performs when transferred against several experiments which distribute total pretraining across a number of networks: two ResNet50 networks each trained for 200k steps, four ResNet50 networks each trained for 100k steps, and five ResNet50 networks each trained for 80k steps.
[…]
All ResNet50 models are pretrained using the MLCommons AlgoPerf training framework and the AdamWScheduleFree optimizer (Kasimbeg et al., 2025), (Defazio et al., 2024). The models are transferred using SGD.
@@ 5.1.2 Results
Our method yield considerable improvements in performance despite maintaining the same total pretraining computation. By combining more networks that have been pretrained for fewer steps, our experiments produce models which generalize significantly better to new datasets: ($+12.5\%$) on iNaturalist18 and ($+6\%$) on CIFAR-100.
As expected, adding more sets of features does not significantly improve performance ImageNet1K. Since ResNet50 models learn features that provide similar information to reduce training loss, combining them on the training data does not does not yield further improvements.
@@
Table 3: Transfer Performance of Concatenated Models, Each Trained For Fewer Iterations A baseline ResNet50 model (cat1) is trained for 450k steps on ImageNet1k. Each experimental method (catn) creates a models by concatenating multiple (n) ResNet50 networks, each trained independently for 400k/n steps. Overall computational effort of pretraining is maintained.
Pretraining Steps Transfer via LinProb

### v3

@@ 4.2 Example - Spurious Features
Size Base-Pairs Score $\uparrow$ Rank $\downarrow$ $\uparrow Imp.\uparrow$
@@
4.3 Example - Genomic Foundation Models
@@
4.4 Example - Rich Representations: The Beginnings of a Solution?
(Zhang and Bottou, 2023) propose learning richer representations to capture these missing features. They concatenating feature representations of individual models and demonstrate significant improvements in accuracy of transfer.
@@ 4.4 Example - Rich Representations: The Beginnings of a Solution?
Table 2: Supervised transfer learning from ImageNet to iNat18, Cifar100, and Cifar10 using linear probing. The cat4 ensemble concatenates representations of 44 separately trained networks. The baseline uses a ResNet model of similar size is trained once. These models contain approximately four times the parameters of the original ResNet50.
@@
[…]
They find the concatenated ensemble to perform comparably to the baseline on ImageNet1K, but far better when transferred to new datasets, ($9\%$ on INat18) and ($5.5\%$ on CIFAR100). Additional comparison to the original ResNet50 reveals that only increasing size of a model (such as the baseline) will worsen performance of transfer: ($5\%$ on INat18), and ($6\%$ on CIFAR100).
@@
This also explains why simply increasing model size without also increasing model richness produces similar or slightly worse results. Observe that these subsets of features work similarly on ImageNet, that is, each set of features provides the same information to reduce training error. Combining these features marginally improves performance on ImageNet1K ($2\%$). Hence, they are not learned during training.
@@
5 Our New Experiments
5.1 Improved Transfer at Fixed Cost
We propose a method to improve transfer performance with fixed total compute. Our method creates rich representations by ensembling, following the method of (Zhang et al., 2022) to concatenate networks pretrained for fewer steps.
In contrast to (Zhang et al., 2022)’s experiments which preserve model parameters (space), our method fixes total compute (time). We call our method (time-concat).
5.1.1 Methodology
Total pretraining time is fixed at 450k iterations per ResNet50 model on 8xH100 GPUs on ImageNet1k (Deng et al., 2009), (He et al., 2016). [^2] Implementation can be found here: https://github.com/facebookresearch/richreps-timecat All pretraining uses the AdamWScheduleFree Optimizer and builds on the MLCommons AlgoPerf Training Framework (Kasimbeg et al., 2025), (Defazio et al., 2024).
  • Baseline Model: A single ResNet50 model fully-trained on ImageNet1k.
  • Ensemble Model: Multiple ResNet50 models trained independently for shorter periods (different seeds, then ensembled).
For example, an ensemble could contain four models trained for $25\%$ of the total time or two models trained for $50\%$ of the total time.
Each ResNet50 network is trained identically except using different random seeds. We transfer these concatenated networks using linear probing on new datasets.
5.1.2 Evaluation
Models are transferred to datasets via linear probing (fine-tuning only the final classification layer) and evaluated on new datasets.
In-Distribution (IID) Performance is measured after transfer to the pretraining dataset (ImageNet1k).
Out-of-Distribution (OOD) Performance is evaluated on datasets not seen during pretraining (iNaturalist18, CIFAR-100, CIFAR-10) (Van Horn et al., 2018), (Krizhevsky, 2009).
Figure 6: Ensembling Boosts ResNet50 Transfer Accuracy Without Extra Compute Baseline: A single (ResNet50Cat1) model is trained for 400k iters on ImageNet1k (plus 5050k iters for a stronger argument). Ensemble: A (ResNet50Cat4) model contains four ResNet50 models trained separately (with different random seeds) on ImageNet for 100k iters each. The ensemble significantly outperforms the baseline during transfer.
22footnotetext: The baseline is pretrained for 450k training steps for a stronger argument.
@@ 5.1.3 Results
Our method yields considerable improvements in performance at fixed training cost.
On ImageNet1k (IID), the ensemble models performed comparably or slightly worse than the single baseline model, showing no difference in their knowledge of ImageNet1k. However, on OOD (Green), the undertrained ensemble significantly outperforms the fully-trained baseline ($12.5\%$ on iNaturalist18, $6\%$ on CIFAR-100).
This performance gap suggests that an ensemble of undertrained models captures patterns relevant to new tasks which a single fully-trained model may miss.
In summary, for a fixed computational budget, an ensemble of intentionally under-trained models outperforms a single, fully-trained model when transferred to unseen distributions.
As expected, concatenating additional sets of features does not significantly boost model performance on ImageNet1K. ResNet50 models generate different sets of features which provide similar information to reduce training loss. Thus, combining them on ImageNet yields no further improvements.
@@
See Table 3 for detailed results.
Table 3: Transfer Performance of Concatenated ResNet models with Fixed Total Compute A baseline ResNet50 model (cat1) is trained for 450k iters on ImageNet1k. Each experimental method (catn) creates a models by concatenating multiple (n) ResNet50 networks, each trained independently for 400k/n iters. Overall computational effort of pretraining is maintained.
Pretraining Iters Transfer via LinProb
@@ Appendix A Additional Empirical Evidence
Appendix B Connections to Multi-Calibration
Missing features fundamentally limits the ability of a model to generalize to new distributions outside the training distribution (OOD). [Wald et al., 2021] finds a link between OOD performance and multi-calibration, introduced by [Hebert-Johnson et al., 2018]. Specifically, they argue that calibration leads to better OOD generalization, by creating invariant representations [Arjovsky et al., 2019]. [Guy et al., 2025] successfully demonstrate this empirically for in-distribution generalization (IID). However, any remediation requires recovering these missing features.